OpenStudy (anonymous):

PLEASE HELP!!April built a model of the F-16 Fighting Falcon that is proportional to the actual F-16 Fighting Falcon. The actual F-16 has a length of 14.8 meters and a wingspan that measures 9.8 meters. Her model has a length of 0.5 meter. What is the approximate measure of the wingspan on April's model?

OpenStudy (anonymous):

.04

OpenStudy (anonymous):

its not .o4

OpenStudy (anonymous):

@deerhunter24 please give an explanation of how you got the answer. Simply giving away the answer might earn you a suspension or warning from the mods.

OpenStudy (anonymous):

oh my b i used the wrong steps

OpenStudy (anonymous):

im tring again

OpenStudy (anonymous):

please explain how youuu got it also so yea i dont know the steps

OpenStudy (anonymous):

i pulled up a 2nd window and got them off of google

OpenStudy (anonymous):

hey meggy will ya fan me

OpenStudy (anonymous):

i will but can you explain how to do it

OpenStudy (anonymous):

yeh one sec let me go back to it

OpenStudy (anonymous):

ok thankss

OpenStudy (anonymous):

take 14.8 and divide it by .5 then take that answer and times it by 9.8 and you should get the correct answer i believe

OpenStudy (anonymous):

@meggy r u there??????????? did u get it or need explanation???????

OpenStudy (anonymous):

let me try what hunter said real fast

OpenStudy (anonymous):

ok when u r over then mention me so that i can help u??????

OpenStudy (anonymous):

ookk well hunter nice try but it waqsnt the correct answer so @Aperogalics :)))wanna give it a shot

OpenStudy (anonymous):

sorry i tried thought it would get you the correct answer sorry

OpenStudy (anonymous):

ohh its ok :) no problem.

OpenStudy (anonymous):

@meggy a model is a copy of that but dimensions are in small ratio like map in map it looks like exact actual but dimensions are made small note that ratio of dimensions of actual to model is same in any direction so length and bredth ratio must be the same so can u evaluate the length ratio??

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

what is the length of actual size plain????????

OpenStudy (anonymous):

sry plane :D

OpenStudy (anonymous):

9.8 meters? :O

OpenStudy (anonymous):

actual F-16 has a length of 14.8 meters got it??????

OpenStudy (anonymous):

ook so the length = 14.8

OpenStudy (anonymous):

ya and it's model has length????????

OpenStudy (anonymous):

.5 meters

OpenStudy (anonymous):

yup:) so ratio is l1/l2=?????????

OpenStudy (anonymous):

.916? how did you get 11/12 im confused

OpenStudy (anonymous):

Don't get confused i m just talking about ratios which must be equal it will be clear wait not 11/12 sry i mean L1/L2=14.8/0.5=29.6 :)

OpenStudy (anonymous):

ook :)

OpenStudy (anonymous):

now their wings must also satisfy this relation:) can you evaluate

OpenStudy (anonymous):

whats that mean?

OpenStudy (anonymous):

simply the ratio must be same every where so wings also has to follow this so actual F-16 has a wingspan that measures 9.8 meters so its model has to have such value that its ratio must be 29.6 so let it be x